# Holographic Aspects of Dynamical Mean-Field Theory

### v1

@@ Abstract
Dynamical mean-field theory (DMFT) is one of the most standard theoretical frameworks for addressing strongly correlated electron systems. Meanwhile, the concept of holography, developed in the field of quantum gravity, provides an intrinsic relationship between quantum many-body systems and space-time geometry. In this study, we demonstrate that these two theories are closely related to each other by shedding light on holographic aspects of DMFT, particularly for electrons with a semicircle density of states. We formulate a holographic renormalization group for the branch Green’s function from the outer edge to the interior of the Bethe lattice network, and then find that its fixed point can be interpreted as a self-consistent solution of Green’s function in DMFT. By introducing an effective two-dimensional anti–de Sitter space, moreover, we clarify that the scaling dimensions for the branch Green’s function and the boundary correlation functions of electrons at the outer edge of the Bethe lattice network are characterized by the fixed-point Green’s function. We also perform DMFT computations for the Bethe-lattice Hubbard model, which illustrate that the scaling dimensions capture the Mott transition in the deep interior.
@@
A key point to discuss the connection between DMFT and holography is that many-body Green’s functions are effectively factorized into local Green’s functions in the limit of infinite spatial dimension. This factorization also implies that DMFT for the semicircle density of states, or equivalently, the Bethe-lattice-type network, captures intrinsic correlation effects in the strongly correlated electron system. The essential feature of the Bethe lattice network is that a loop-free network structure and translational invariance coexist at the level of an infinite-size tree network, and the number of nodes increases exponentially toward the outer edge (see Fig. 1). Interestingly, the Bethe lattice network is also known as a Burhat-Tits tree in the context of $p$-adic AdS/CFT [42, 43, 44, 45], relevant to the holographic RG for the Ising model [46]. These features prompt us to set up a holographic RG transformation for the branch Green’s function by taking the partial sum from the outer edge to the interior of a correlated electron system on the Bethe lattice network. Then, we explain that the bulk fixed point equation of the holographic RG corresponds to the self-consistent relation in DMFT. By appropriately smearing out lattice nodes in the network, one can introduce an effective coordinate corresponding to Poincaré’s upper half [46]. We then demonstrate that the scaling dimensions for the correlation functions of electrons on the outer edge boundary can be related to the convergence of the holographic RG iteration in DMFT, in consistency with the $p$-adic AdS/CFT description. Here, we would like to emphasize that, in contrast to the Bethe lattice Ising model, the holographic formulation for DMFT incorporates quantum effects through the imaginary-time path integral.
@@ II.5 Boundary correlation functions
In a similar way, we can also consider the correlation function of the composite operator, which is relevant to the holographic argument in the next section. Taking into account Eq. (53), we analyze the correlation function of the normal ordered operators
@@ IV Numerical results
We have reformulated dynamical mean-field theory (DMFT) for the semicircle density of states as a holographic renormalization group (RG). In particular, we recursively determine the branch Green’s function of electrons from the outer edge into the interior of the corresponding Bethe lattice network, where the fixed point relation of the recursive RG transformation yields the self-consistent solution of the local Green’s function in DMFT. Moreover, we have introduced an effective AdS2 with the Poincaré radius $L=1/\log p$ near the outer edge boundary, by smearing out the lattice nodes in the Bethe lattice. We have then found that the scaling dimensions for the branch Green’s function and the correlation function of the boundary electron density operators satisfy the scalar field relation (77) in AdS2. This relation of the scaling dimensions is also consistent with $p$-adic AdS/CFT. We have performed DMFT calculations for the Hubbard model with a semicircle density of states; the scaling dimensions corresponding to the metallic and insulating solutions coexist in $2.36<U<2.55$, reflecting that the Mott transition at a finite temperature is of first order. We have demonstrated that the scaling dimensions for the excitations detect the features of the metallic and insulating phases, implying that the boundary correlation functions can systematically capture information about the deep interior. In this paper, we have restricted ourselves to the paramagnetic case. However, the Bethe lattice has a bipartite structure, implying that the ground state may have an antiferromagnetic order[31, 32, 34, 58] An extension of the holographic RG for the magnetically ordered system is an interesting problem.
@@
The holographic RG formulation of DMFT provides several interesting implications associated with the holography and quantum entanglements. In the following, we briefly discuss some issues in order. First, in this work, we have concentrated on the Bethe lattice model, where we could directly introduce the recursive relation of Green’s function. For general lattice models, however, it is not straightforward to write down the recursive relation directly on the real-space lattice network. In DMFT, instead, the local self-energy is extracted from Green’s function, and then the lattice network information is self-consistently implemented into the local Green’s function through the Hilbert transformation with respect to the density of states, where the $k$ dependence of the single-particle dispersion is traced out. The use of the Hilbert transformation suggests that the real-space network structure could be appropriately smeared out without directly referring to the real-space lattice network. How the DMFT framework generates a curved space coordinate $(r,x)$ relevant to AdS2 would be an important problem.

### v2

@@ Abstract
Dynamical mean-field theory (DMFT) is one of the most standard theoretical frameworks for addressing strongly correlated electron systems. Meanwhile, the concept of holography, developed in the field of quantum gravity, provides an intrinsic relationship between quantum many-body systems and space-time geometry. In this study, we demonstrate that these two theories are closely related to each other by shedding light on holographic aspects of DMFT, particularly for electrons with a semicircle density of states. We formulate a holographic renormalization group for the branch Green’s function from the outer edge to the interior of the Bethe lattice network, and then find that its fixed point can be interpreted as a self-consistent solution of Green’s function in DMFT. By introducing an effective two-dimensional anti-de Sitter space, moreover, we clarify that the scaling dimensions for the branch Green’s function and the boundary correlation functions of electrons at the outer edge of the Bethe lattice network are characterized by the fixed-point Green’s function. We also perform DMFT computations for the Bethe-lattice Hubbard model, which illustrate that the scaling dimensions capture the Mott transition in the deep interior.
@@
A key point in discussing the connection between DMFT and holography is that many-body Green’s functions are effectively factorized into local Green’s functions in the limit of infinite spatial dimension. This factorization also implies that DMFT for the semicircle density of states, or equivalently, the Bethe-lattice-type network, captures intrinsic correlation effects in the strongly correlated electron system. Then, an essential feature of the Bethe lattice network is that a loop-free network structure and translational invariance coexist at the level of an infinite-size tree network, and the number of nodes increases exponentially toward the outer edge (see Fig. 1), which allows us to interpret the Bethe lattice Ising model as a toy model in the gravity side of Euclidean AdS2 [42], consistent with a Burhat-Tits tree in the context of $p$-adic AdS/CFT [43, 44, 45, 46].
Inspired by the observations above, we set up a holographic RG transformation by taking a partial sum from the outer edge to the interior of a correlated-electron system on a Bethe lattice network. Assuming the effective factorization of Green’s function for a descendant branch in the network as in DMFT, we derive a holographic RG relation for the branch Green’s function, which corresponds to the iterative update of an effective chemical potential in a DMFT calculation. We then find that the bulk fixed point equation of the holographic RG corresponds to the self-consistent relation in DMFT. Moreover, we can calculate the correlation function of the electron-density operators, which involves the response of the system to a perturbation of the chemical potential at the outer edge node. By appropriately smearing out lattice nodes in the network, moreover, one can also introduce an effective coordinate corresponding to Poincaré’s upper half [42]. We then extract the scaling dimensions associated with the outer-edge chemical potential, which can be related to the convergence of the DMFT iteration, suggesting that the DMFT for the correlated-electron system can also serve as a prototypical lattice-network model on the gravity side that incorporates quantum effects via the imaginary-time path integral.
@@ II.2 Fourier representation
Here, we discuss the physical interpretation of Eq. (27). By introducing an effective chemical potential with
$\tilde{\mu}_{N-\bar{n}}(i\omega_{l})\equiv\mu-pt^{2}G_{N-\bar{n}+1\sigma}(i\omega_{l})\,$, (29)
we can regard Eq. (27) as an RG transformation for the $\omega$- (or equivalently $\tau$-) dependent chemical potential. Then, Eq. (27) describes how the chemical potential applied at the outer edge node is renormalized toward the interior: $\tilde{\mu}_{N}\to\tilde{\mu}_{N-1}\to\tilde{\mu}_{N-2}\to\cdots$. In other words, the iterative update of ${G}(i\omega_{l})$ plays the same role as a holographic RG iteration of the effective chemical potential, as in the case of the magnetic field for the Bethe lattice Ising model.
@@ II.3 Fixed point and linearized RG equation
Here, note that the linearized RG equation (35) is written as $\delta\tilde{\mu}_{N-\bar{n}}(z)=\lambda_{\sigma}(z)\delta\tilde{\mu}_{N-\bar{n}+1}(z)$ in terms of the renormalization of the effective chemical potential.
@@ II.5 Boundary correlation functions
In a similar way, we can also consider the correlation function of the composite operator, which is relevant to the holographic argument in the context of the response function against a perturbation with respect to $\tilde{\mu}(\omega)$ in the next section. Taking into account Eq. (54), we analyze the correlation function of the normal ordered operators
@@ IV Numerical results
We have reformulated dynamical mean-field theory (DMFT) for the semicircle density of states as a holographic renormalization group (RG). In particular, we recursively determine the branch Green’s function of electrons from the outer edge into the interior of the corresponding Bethe lattice network, where the fixed point relation of the recursive RG transformation yields the self-consistent solution of the local Green’s function in DMFT. Moreover, we have introduced an effective AdS2 with the Poincaré radius $L=1/\log p$ near the outer edge boundary, by smearing out the lattice nodes in the Bethe lattice. We have then found that the scaling dimensions for the branch Green’s function (or equivalently for the effective chemical potential) and the correlation function of the boundary electron density operators satisfy the scalar field relation (78) in AdS2. This relation of the scaling dimensions is also consistent with $p$-adic AdS/CFT. We have then performed DMFT calculations for the Hubbard model with a semicircle density of states; the scaling dimensions corresponding to the metallic and insulating solutions coexist in $2.36<U<2.55$, reflecting that the Mott transition at a finite temperature is of first order. We have demonstrated that the scaling dimensions for the excitations detect the features of the metallic and insulating phases, implying that the boundary correlation functions can systematically capture information about the deep interior. In this paper, we have restricted ourselves to the paramagnetic case. However, the Bethe lattice has a bipartite structure, implying that the ground state may have an antiferromagnetic order[31, 32, 34, 58] An extension of our holographic RG for the magnetically ordered system is an interesting problem.
In the context of AdS/CFT, the above results suggest that, reflecting the particular nature of the Bethe lattice, DMFT can serve as a prototypical toy model that incorporates quantum fluctuations on the bulk gravity side. On the other hand, we have not directly discussed the duality with the corresponding boundary CFT, which seems highly nontrivial, since the outer edge nodes of the Bethe lattice exhibit a highly fragmented structure in the circumference direction. In this sense, our approach to DMFT should be distinct from the bulk reconstruction based on the boundary CFT[60]. Such a situation could be termed “AdS/DMFT”. Understanding the nature of the boundary theory corresponding to DMFT is an interesting future problem.
@@
In addition, the holographic RG formulation of DMFT provides several interesting implications associated with the holography and quantum entanglements. In the following, we briefly discuss some issues in order. First, in this work, we have concentrated on the Bethe lattice model, where we could directly introduce the recursive relation of Green’s function. For general lattice models, however, it is not straightforward to write down the recursive relation directly on the real-space lattice network. In DMFT, instead, the local self-energy is extracted from Green’s function, and then the lattice network information is self-consistently implemented into the local Green’s function through the Hilbert transformation with respect to the density of states, where the $k$ dependence of the single-particle dispersion is traced out. The use of the Hilbert transformation suggests that the real-space network structure could be appropriately smeared out without directly referring to the real-space lattice network. How the DMFT framework generates a curved space coordinate $(r,x)$ relevant to AdS2 would be an important problem.